OpenStudy (anonymous):

What is the solution to the equation log4 x + log4 8 = 3 ? x = one–half x = 2 x = 8 x = 32

OpenStudy (anonymous):

I'd like to know whether 4 is the base or just being multiplied ?

OpenStudy (anonymous):

i think its the base

OpenStudy (anonymous):

the base

OpenStudy (anonymous):

Using the log property\[\log_{4}(8x) = 3 \]\[8x = 4^{3}\]\[x = \frac{ 64 }{ 8 } = 8\]

OpenStudy (anonymous):

\[\log_{4}( x)+ \log_{4}(8)= \log_{4}(8x)=3, 4^{3}=(8x), 64=(8x), x=64/8=8\]